Angle bisectors of angles BAC and BCA of triangle ABC intersect its altitude BH at points P and
Q respectively. It is known that point H lies on side AC, P is situated between H and Q,
BQ = 54,5, QP = 17,5, PH = 36. Find the area S of triangle ABC. As the answer,
indicate the value of S√3 .
1
Expert's answer
2020-11-08T18:42:52-0500
"BH=BQ+QP+PH="

"=54.5+17.5+36=108"

The right triangle "APH"

Let "AH=x, \angle PAH=\alpha." Then "\tan(\alpha)=\dfrac{PH}{AH}=\dfrac{36}{x}"

The right triangle "ABH"

"\tan(2\alpha)=\dfrac{BH}{AH}=\dfrac{108}{x}"

"\tan(2\alpha)=\dfrac{2\tan(\alpha)}{1-\tan^2(\alpha)}=\dfrac{108}{x}"

"\dfrac{108}{x}=\dfrac{2\cdot\dfrac{36}{x}}{1-(\dfrac{36}{x})^2 }""1-(\dfrac{36}{x})^2=\dfrac{2}{3}"

"x=36\sqrt{3}"

"\alpha=30\degree"

The right triangle "QCH"

Let "CH=y, \angle QCH=\beta." Then "\tan(\beta)=\dfrac{QH}{CH}=\dfrac{53.5}{y}"

The right triangle "BCH"

"\tan(2\beta)=\dfrac{BH}{CH}=\dfrac{108}{y}"

"\tan(2\beta)=\dfrac{2\tan(\beta)}{1-\tan^2(\beta)}=\dfrac{108}{y}"

"\dfrac{108}{y}=\dfrac{2\cdot\dfrac{53.5}{y}}{1-(\dfrac{53.5}{y})^2 }""1-(\dfrac{53.5}{y})^2=\dfrac{107}{108}"

"(\dfrac{y}{53.5})^2=108"

"y=321\sqrt{3}"

Then

"AC=AH+CH=36\sqrt{3}+321\sqrt{3}=357\sqrt{3}"

"S_{ABC}=\dfrac{1}{2}BH\cdot AC="

"=\dfrac{1}{2}(108)(357\sqrt{3})=19278\sqrt{3}"

"S\sqrt{3}=57834"